From the sum of 4 + 3x and 5 – 4x + 2x 2 , subtract the sum of 3x 2 – 5x and –x 2 + 2x + 5.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

First we have to find out the sum of 4 + 3x and 5–4x + 2x^2
= 4 + 3x + (5–4x + 2x^2)
= 4 + 3x + 5–4x + 2x^2
= 4 + 5 + 3x–4x + 2x^2
= 9 - x + 2x^2
= 2x^2–x + 9... [equation 1]
Then, we have to find out the sum of 3x^2–5x and–x^2+ 2x + 5
= 3x^2–5x + (-x^2+ 2x + 5)
= 3x^2–5x–x^2+ 2x + 5
= 3x^2–x^2–5x + 2x + 5
= 2x^2–3x + 5... [equation 2]
Now, we have to subtract equation (2) from equation (1)
= 2x^2- x + 9–(2x^2–3x + 5)
= 2x^2–x + 9–2x^2+ 3x–5
= 2x^2–2x^2–x + 3x + 9–5
= 2x + 4

​Thanks